OpenStudy (anonymous):

Simplify the expression and show your work x²-6+9 x²-2x-3 ------- ÷ ------- 2x³+5x 4x²+10

OpenStudy (anonymous):

dividing by a fraction is the same as multiplying by its reciprocal

OpenStudy (anonymous):

and how do you do that?

OpenStudy (anonymous):

example \[\frac{a}{b} \div \frac{c}{d} = \frac{a}{b}* \frac{d}{c}\]

OpenStudy (anonymous):

oh so i multiply?

OpenStudy (anonymous):

multiply by the reciprocal

OpenStudy (anonymous):

notice how the d and c switched places